# arXiv:2603.01972v1[eess.SY] 02 Mar 2026

### v1

@@ 2 Common Pitfalls to Convergence
Pitfall 2: Fragmented ontologies and sectoral silos. Most tools are developed within individual sectors–electricity, water, transportation, food–without recognizing the interconnected systems they influence or depend upon [Farid:2016:ISC-BC06, Farid:2022:ISC-J49, Farid:2022:01]. This fragmentation leads to ontological mismatches when attempting integration [monteiro:2009:01]. Researchers attempting multi-sector modeling must reconcile conflicting problems, definitions, data structures, and modeling assumptions across disciplines [iyer:2019:00, reed:2022:00, srikrishnan:2022:00]. For instance, national decarbonization scenarios often integrate detailed power sector models with coarse or static representations of water and land systems, leading to unrealistic assumptions about cooling water availability or biofuel feedstock supply [kraucunas:2015:00]. Even ambitious efforts like food-energy-water nexus studies often fall short: a systematic review of 245 publications revealed that most did not actually capture all three dimensions, let alone their interactions [Albrecht:2018:00]. As the Intergovernmental Panel on Climate Change (IPCC) 2023 synthesis report states: “Maladaptation can be avoided by flexible, multi-sectoral, inclusive, long-term planning…” [ipcc:2023:00]. Without common ontologies and standards, attempts at integration often result in incoherent or incompatible models [holzinger:2005:00, razavi:2025:00].
@@
Pitfall 4: Co-simulation limitations in structure and scale. Co-simulation is a popular technique for connecting models of different systems, but it usually assumes a sequential or looped flow of information (i.e., input/output pairs)[gomes:2017:00]. This structure is ill-suited for modeling multilateral, bi-directional SoS with complex feedbacks and distributed agency [Taveres-Cachat:2021:00]. Moreover, co-simulation methods are rarely equipped to manage the emergence of disparate spatial and temporal scales that arise in climate systems, infrastructure transitions, and social dynamics [barbierato:2022:00, alfalouji:2023:00]. For example, co-simulating urban energy, water, and transportation networks has proven difficult because decisions about land-use or infrastructure investment unfold over decades, whereas operational control models run on hourly timescales [turner:2017:00, van-dam:2012:00]. In the context of accelerating global change, we need tools that allow for multi-timescale emergence, with the ability to represent slow-evolving institutional transformations alongside rapid technological shocks [lemke:2000:00]. Without this, convergence modeling collapses under either oversimplification or unmanageable complexity.
@@
Pitfall 6: Not communicating the benefits of convergence Collectively, these five pitfalls reveal not isolated shortcomings, but a shared structural deficiency in the way science organizes knowledge and practice. They expose the absence of a unifying scientific and institutional infrastructure capable of linking discipline-specific tools, data, and governance systems into a coherent whole. Without such integration, existing approaches will continue to generate fragmented, mismatched, or non-transferable solutions, mirroring policy silos in which government agencies pursue parallel mandates while overlooking critical cross-sector dependencies [Cumming:2013:00, Morrison:2017:00].
@@ 3.1 Knowledge Generation Described by a Meta-Cognition Map
The Visual Domain. The visual domain plays a pivotal role in bridging informal thought and formal representation [jankun-kelly:2007:00]. Every discipline has developed its own visual abstractions, including free-body diagrams in mechanics, circuit diagrams in electrical engineering, process flow diagrams in chemical engineering, and GIS maps in geospatial disciplines. These tools enable communication, iteration, and collaborative model building [valencia-garcia_technologies_2021, al-kodmany:2001:00]. However, they are deeply discipline-specific. This specialization can become a barrier when collaborating across domains, particularly in SoS contexts. Cross-disciplinary collaboration demands that visual tools be ontologically coherent and extensible across systems [aagaardhansen:2007:00]. SysML, for instance, offers a structured visual syntax that can span mechanical, cyber, and environmental systems [mann:2009:00].
@@ 4 Case Study: Demonstration of the SoS Convergence Paradigm in the Chesapeake Bay Watershed
(b) Connected watershed [harris:2025:01]
@@
From a convergence perspective, the Chesapeake Bay Watershed concentrates the hallmark challenges of Anthropocene systems: scientific complexity, institutional fragmentation, and socio-economic coupling [Morrison:2017:00, Hood:2021:00]. The CBP’s modeling system and the Chesapeake Assessment Scenario Tool (CAST) provide valuable data and calibration, yet remain structurally opaque and difficult to extend beyond domain boundaries [Hood:2021:00]. In the CBP context, opacity manifests in routine yet undocumented practices like unofficial spreadsheet exchanges, previewed model runs and negotiated resubmissions, use of state/consultant alternatives (e.g., VAST/MAST), and consensus-based deferral of QA/QC which the dual-layer overlay makes explicit [jangjoo:2025:00]. These conditions call for a framework that satisfies the six design criteria of convergence–systems foundation, common ontology, extensibility, analytical and synthetic capacity, and problem independence–while facilitating translation across the five cognitive domains of the meta-cognition map: Real-World, Systems Thinking, Visual, Mathematics, and Computing.
@@ 4.1 SoS Computational Framework
This approach was then extended to the full Chesapeake Bay Watershed (Fig. 4). A SysML–HFGT instantiation now captures the hierarchical structure of elements of the watershed, including land segments, outlet points, river networks, and estuarine boundaries [harris:2025:01]. Geospatial features from CAST were translated into tensor-based HFGT representations that were then used in coordination with CAST behavioral data to simulate the movement of nitrogen and phosphorus throughout the watershed (Fig. 4(b)). The project utilizes the Weighted Least Squares Error Hetero-functional Graph State Estimator (WLSE–HFGSE) introduced by Thompson et al. [Thompson:2025:ISC-JR11] to overcome the uncertainty inherent in large, data-sparse systems such as in the CAST model . When adapted to a watershed system, this estimator infers unobserved water and nutrient flows from partial data while maintaining physical consistency [harris:2025:01]. The result was a transparent, extensible model that mirrors CAST’s routing assumptions while remaining fully aligned with MBSE principles. Rather than replicating CAST, this ongoing modeling effort seeks to augment it, revealing structural relationships withing the watershed system and establishing a foundation for future extension to additional environmental or socio-technical systems.
In parallel, Naderi et al. [naderi:2025:01] demonstrated the methodological generality of the SysML–HFGT framework by applying it to economic input–output (I–O) models. Using a Rectangular Choice of Technology (RCOT) example [Duchin:2011:00], they showed that conventional I–O dynamics can be reproduced within HFGT’s mathematical formalism, providing structural clarity and enabling integration with environmental systems. This offers an initial application of the convergence framework to economic process modeling for future extension to the Chesapeake Bay Watershed region to coordinate with hydrological processes.

### v2

# A System-of-Systems Convergence Paradigm for Societal Challenges of the Anthropocene

@@ 2 Common Pitfalls to Convergence
Pitfall 2: Fragmented ontologies and sectoral silos. Most tools are developed within individual sectors–electricity, water, transportation, food–without recognizing the interconnected systems they influence or depend upon [Farid:2016:ISC-BC06, Farid:2022:ISC-J49, Farid:2022:01]. This fragmentation leads to ontological mismatches when attempting integration [monteiro:2009:01]. Researchers attempting multi-sector modeling must reconcile conflicting problems, definitions, data structures, and modeling assumptions across disciplines [iyer:2019:00, reed:2022:01, srikrishnan:2022:00]. For instance, national decarbonization scenarios often integrate detailed power sector models with coarse or static representations of water and land systems, leading to unrealistic assumptions about cooling water availability or biofuel feedstock supply [kraucunas:2015:00]. Even ambitious efforts like food-energy-water nexus studies often fall short: a systematic review of 245 publications revealed that most did not actually capture all three dimensions, let alone their interactions [Albrecht:2018:00]. As the Intergovernmental Panel on Climate Change (IPCC) 2023 synthesis report states: “Maladaptation can be avoided by flexible, multi-sectoral, inclusive, long-term planning…” [ipcc:2023:00]. Without common ontologies and standards, attempts at integration often result in incoherent or incompatible models [holzinger:2005:00, razavi:2025:00].
@@
Pitfall 4: Co-simulation limitations in structure and scale. Co-simulation is a popular technique for connecting models of different systems, but it usually assumes a sequential or looped flow of information (i.e., input/output pairs)[gomes:2017:00]. This structure is ill-suited for modeling multilateral, bi-directional SoS with complex feedbacks and distributed agency [taveres-Cachat:2021:01]. Moreover, co-simulation methods are rarely equipped to manage the emergence of disparate spatial and temporal scales that arise in climate systems, infrastructure transitions, and social dynamics [barbierato:2022:00, alfalouji:2023:00]. For example, co-simulating urban energy, water, and transportation networks has proven difficult because decisions about land-use or infrastructure investment unfold over decades, whereas operational control models run on hourly timescales [turner:2017:00, van-dam:2012:00]. In the context of accelerating global change, we need tools that allow for multi-timescale emergence, with the ability to represent slow-evolving institutional transformations alongside rapid technological shocks [lemke:2000:00]. Without this, convergence modeling collapses under either oversimplification or unmanageable complexity.
@@
Pitfall 6: Not communicating the benefits of convergence Collectively, these five pitfalls reveal not isolated shortcomings, but a shared structural deficiency in the way science organizes knowledge and practice. They expose the absence of a unifying scientific and institutional infrastructure capable of linking discipline-specific tools, data, and governance systems into a coherent whole. Without such integration, existing approaches will continue to generate fragmented, mismatched, or non-transferable solutions, mirroring policy silos in which government agencies pursue parallel mandates while overlooking critical cross-sector dependencies [Cumming:2013:00, morrison:2017:01].
@@ 3.1 Knowledge Generation Described by a Meta-Cognition Map
The Visual Domain. The visual domain plays a pivotal role in bridging informal thought and formal representation [jankun-kelly:2007:00]. Every discipline has developed its own visual abstractions, including free-body diagrams in mechanics, circuit diagrams in electrical engineering, process flow diagrams in chemical engineering, and GIS maps in geospatial disciplines. These tools enable communication, iteration, and collaborative model building [valencia-garcia:2021:00, al-kodmany:2001:00]. However, they are deeply discipline-specific. This specialization can become a barrier when collaborating across domains, particularly in SoS contexts. Cross-disciplinary collaboration demands that visual tools be ontologically coherent and extensible across systems [aagaardhansen:2007:00]. SysML, for instance, offers a structured visual syntax that can span mechanical, cyber, and environmental systems [mann:2009:00].
@@ 4 Case Study: Demonstration of the SoS Convergence Paradigm in the Chesapeake Bay Watershed
(b) Connected watershed [harris:2026:00]
@@
From a convergence perspective, the Chesapeake Bay Watershed concentrates the hallmark challenges of Anthropocene systems: scientific complexity, institutional fragmentation, and socio-economic coupling [morrison:2017:01, Hood:2021:00]. The CBP’s modeling system and the Chesapeake Assessment Scenario Tool (CAST) provide valuable data and calibration, yet remain structurally opaque and difficult to extend beyond domain boundaries [Hood:2021:00]. In the CBP context, opacity manifests in routine yet undocumented practices like unofficial spreadsheet exchanges, previewed model runs and negotiated resubmissions, use of state/consultant alternatives (e.g., VAST/MAST), and consensus-based deferral of QA/QC which the dual-layer overlay makes explicit [jangjoo:2025:00]. These conditions call for a framework that satisfies the six design criteria of convergence–systems foundation, common ontology, extensibility, analytical and synthetic capacity, and problem independence–while facilitating translation across the five cognitive domains of the meta-cognition map: Real-World, Systems Thinking, Visual, Mathematics, and Computing.
@@ 4.1 SoS Computational Framework
This approach was then extended to the full Chesapeake Bay Watershed (Fig. 4). A SysML–HFGT instantiation now captures the hierarchical structure of elements of the watershed, including land segments, outlet points, river networks, and estuarine boundaries [harris:2026:00]. Geospatial features from CAST were translated into tensor-based HFGT representations that were then used in coordination with CAST behavioral data to simulate the movement of nitrogen and phosphorus throughout the watershed (Fig. 4(b)). The project utilizes the Weighted Least Squares Error Hetero-functional Graph State Estimator (WLSE–HFGSE) introduced by Thompson et al. [Thompson:2025:ISC-JR11] to overcome the uncertainty inherent in large, data-sparse systems such as in the CAST model . When adapted to a watershed system, this estimator infers unobserved water and nutrient flows from partial data while maintaining physical consistency [harris:2026:00]. The result was a transparent, extensible model that mirrors CAST’s routing assumptions while remaining fully aligned with MBSE principles. Rather than replicating CAST, this ongoing modeling effort seeks to augment it, revealing structural relationships withing the watershed system and establishing a foundation for future extension to additional environmental or socio-technical systems.
In parallel, Naderi et al. [naderi:2026:00] demonstrated the methodological generality of the SysML–HFGT framework by applying it to economic input–output (I–O) models. Using a Rectangular Choice of Technology (RCOT) example [Duchin:2011:00], they showed that conventional I–O dynamics can be reproduced within HFGT’s mathematical formalism, providing structural clarity and enabling integration with environmental systems. This offers an initial application of the convergence framework to economic process modeling for future extension to the Chesapeake Bay Watershed region to coordinate with hydrological processes.